FEL Upgrade Project Weekly Brief –August 5-9, 2002




Three more GWs were received from the vendor which completes the group required for the IR machine. They will be characterized in the magnet test stand before installation. Virtually all of the remaining stands were installed in the tunnel and substantial progress is now being made on installation now that full crews have returned to the FEL from the main machine maintenance period.

One of the test/assembly stands was set up with a complete deformable mirror assembly, actuators and LVDT. This will permit testing of the alignment accuracy and response frequency with the full mass driven.


Preparations and dry runs were held for the DOE Director of the Office of Science,

Ray Orbach's, visit to JLab on August 19 and 20.

On August 8 we briefed Mr. Peter Teets, the Under SECAF(Space/Director, NRO) and his Deputy for Military Space, Robert Dickman on the Navy scale-up program.

We thank the Office of Naval Research for sending us an additional $80k of the $300k remaining FY02 allocation for the agreed tasks.

WBS 3 (Beam Physics):

A pi-mode based n-cell cavity model for user-defined RF wavelength was constructed. Beamlette stacking was investigated.

WBS 4 (Injector):

The SF6 aluminum tank is ready for shipment to be modified. Bids are expected by Tuesday next week. The fixtures for welding the support tube to its mounting flange are being fabricated in the machine shop.

Continued to work with W&M on the implantation of the support tube for the ball cathode. Two more runs free of arcs were completed this week, finishing the implantation on the end where the ball is attached. The first of these runs was successfully performed at 15 kV. However, instabilities in the pulse current were again present in the second run, forcing to implant at only 2 kV. The fifth run on the opposite end is in progress today at full 15 kV. The next and final run will be performed on Tuesday next week.

Gun HVPS - The upper sections of the Driver and Measurement stacks were installed this week. The middle toroids are in the Machine Shop being drilled for assembly next week. The pipe for the relief valve was also fabricated.

WBS 5 (SRF):

Studies of HOM damping in the 7 cell cavities continue including Q measurements on the actual cavities from the first (CEBAF) module and calculations of stability limits using TDBBU and other codes. All 8 cavities of this first CEBAF module are to be passed over for string assembly next week.

WBS 6 (RF):

Quarter HVPS - Continued debugging the hardware and P/C boards. Initial talks were held with G. Lahti to get started on the communication link between the PLC in the HVPS and EPICS for remote operation and display. The waveguide is being connected for unit 3 to enable No-Load testing of the HVPS early next week. The interlocks have all been checked.

Quarter Klystrons - The filament resistance of the 100 kW klystrons appear to be slowly increasing. A. Mizuhara is not concerned at this time. We will continue to monitor the klystron now that the instrumentation is settling down.

Zone 4 - The missing waveguide components are on order or at the Machine Shop for fabrication.

WBS 8 (Instrumentation):

Drive Laser Pulse Control (DLPC) - A 64bit digital I/O card (GPVS64) has been installed into iocfel10. This board will provide the necessary control for the CO305 card (for micropulse control) and the basic user requests to open/insert and close/remove the MPS shutter and the ND2 filter respectively. New software screens for macropulse control have been created and will be test soon.

Beam Loss Monitors (BLMs) - The fabrication drawings for the blm front board (F0151) have been signed-off and the board has been sent out for prototype production with a 5-day turn. The majority of the parts are in-house and ready for assembly. The assembly and testing will be done in stages. The BLM rear board (F0150) is being ECO'ed to correct an component placement issue in regards to mating with the cable from the BLM heads. Additionally, the requirements for the software have been worked out as well and will be ready as the F0151board goes into its testing phase.

Completed the cable installation for the Magnet Personel Safety Interlock. Terminations and checkout are still in progress. The PSS RF interface cables have been terminated into the PSS PLC's and into their interface chassis in Zone 4. Work has been completed on the assembly of the super trim racks in Zones 1 and 5. Klixon cables have been pulled into their new destinations. The BPM cables have dressed in and labeled in Zone 4.

The DSP code is progressing for the spectrometer (and to be used for all other apps. Like BPM, power supplies...). This weeks effort by our student was on running the routine for outputting a waveform (serves to form the pulse) from a GPIO pin when a signal is fed into one of the timer pins (serves to form the beam sync). The signal received from the GPIO is a bit noisy (either a glitch in the code or needs filtering, hence executing code that would remove the noisy component from the signal(this will prove helpful in the future as well).

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

We received the next three GWs.

• The remaining 6 cores of the second batch of 6 Magnets for UV are undergoing gap checking.

Injector Dipoles (DU/DV)

• The first GV started magnet measurements on the magnet test stand.

• The first attempt at gluing shim materials to GU cores was not correct. They are doing more tests.

• The remaining 5 GVs are ready to ship early next week.

Arc 180 Degree Dipoles (GY)

• At Wang NMR, the winding of the fourth GY Coil is complete.

• GY Coil Potting fixture remains in fabrication.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

• Process Equipment Co. (PECo) will ship the four GQ cores on Monday. The GX cores are on track for shipping in two weeks.

• At Wang NMR, potting fixtures remain in fabrication for the GQ and GX coils.


Trim Quad (QT)

• The first unit was tested using the movable traverse method on the quad test stand. The data is being analyzed.

Sextupole (SF)

• DULY Research is finishing the Sextupole drawings per our mark-ups.

Octupole (OT)

• DULY continued design of the Octupole.

Beam Line and Vacuum

• For the Arc Chamber job, Master Machine is awaiting samples from the 5-axis water jet cutting vendor before authorizing the cutting the chamber parts. Remaining parts are in machining.

• The Jlab Shop is starting the full penetration welding of the X Chamber while the end transitions pieces (being made by Master Machine) are nearly ready for attachment.

• The Jlab Shop is starting fabrication of the four chambers for the optical chicane region.

• Design of the regions around the wiggler is 80 % complete.

• The stands for sextupoles and trim quads are in fabrication.

• Design of the Octupole stand continued.

• Girder assembly has started up again with the return of technicians from their CEBAF duties during the latest down period. This time, corrector dipoles which have been characterized and signed off by David Douglas, will be added to the assembly process.

• We issued another revision (L) of the power supply list that included the new requirement for the 4 GC vertical corrector dipoles in the Arcs which will utilize the steel and some of the windings of four QT Trim Quadrupoles. For each of the four systems, a standard trim card power supply is replaced by two low current (2A), power supplies that float with respect to ground and are able to be controlled accurately while run in parallel.

WBS 11 (Optics):

This week's work on the optical cavity we implemented a fix to reduce the binding problems encountered with the Si:bronze/sputtered-silver coated leadscrew. This ameliorated the problem, but did not eliminate it. We did find a component that wasn't made to our drawing tolerances, and fixing that will reduce the problem. Even with the binding problem, the torque of this nut/leadscrew combination was lower (~ 22 in-lbs) than the Vespel/bare stainless steel leadscrew combination.

One of the test/assembly stands was set up with a complete deformable mirror assembly, actuators and LVDT. The assembly went fairly smoothly. We are using this and a portion of the OCMMS test stand to look for low frequency resonances that could prove troublesome when we install the optical cavity hardware in the FEL vault. The good news is that no mechanical resonances attributable to the cavity hardware have been found. We are using this setup to continue working on the OCMMS feedback circuits. We operated a prototype of the utility (water , power and instrumentation) routing mechanism in excess of 2500 times to see if there were any problems. Some binding and resultant particulate generation was noted, an will be addressed. The contract for the cavity vacuum vessels was awarded. We anticipate receipt near the end (27th) of September.

We measured the phase noise of our frequency-locked ultrafast laser. When everything is optimized, the value corresponds to a timing jitter of about 400 fs rms. However, there are times when the timing jitter is on the order of 1 ps; too high in many cases to do the measurements we would like to do. We will take more data next week and see if we can determine if this behavior was a fluke, or a problem with the vendor's locking implementation.